a) Assume KBL Limited and Richet International are multinational firms that are weighing-in the option of simultaneously entering the Ghanaian market for the first time. If neither enters, both earn a payoff of zero. If both enter, they both lose 300. If one firm enters, it gains 150 while the other earns zero. i)Set up the payoff matrix for this game and determine if any Nash equilibria exist. (5 marks) ii)Can you predict the outcome of this game? (3marks) iii) What is the outcome of game if KBL gets to decide first? Explain you answer (5 marks)

1
Expert's answer
2021-11-11T18:14:20-0500

i) The payoff matrix for this game is:



There exists a Nash equilibrium in mixed strategies.

ii) Outcome of this game:

"\frac{-300}{-600} + \frac{150}{600}\\
\frac{1}{2} + \frac{1}{4}\\
0.5 + 0.25\\
= 0.75"


ii)If the KBL decides first, the outcome will be "\frac{-300}{-600} = \frac{1}{2}" , which is the outcome of Richet International not entering the market.

This is because Richet International will not enter into the market early, making the KBL Limited to enter into the market and earn a profit of 150